Two marbles are drawn randomly one after the other without replacement from a jar that contains 4 red marbles, 2 white marbles, and 7 yellow marbles. Find the probability of the following events.


(a) A red marble is drawn first followed by a white marble.
The probability is :


(b) A white marble is drawn first followed by a white marble.
The probability is :


(c) A yellow marble is not drawn at all.
The probability is :
Answer #1

Ans:

a)P(first is red and second is white)=(4/13)*(2/12)=8/156

b)P(first is white and second is white)=(2/13)*(1/12)=2/156

c)P(not yellow)=(6/13)*(5/12)=30/156=15/78